OpenStudy (anonymous):

Suppose that x is an acute angle of a right triangle and that tan(x)=(3 square root 2)/(8). Find sec(x) and csc (x).

OpenStudy (anonymous):

Tan is opp/ adj therefore, opposite = 3root2 and adj= 8. To find the hyp. use a^2 +b^2 = c^2 to find that hyp = root70 You're looking for sec which is hyp/adj or root70/8 and csc which is hyp/opp or root70/3root2

OpenStudy (anonymous):



OpenStudy (anonymous):

find the hypotenuse using pythagoras. then use the trig ratios secant is hypotenuse over adjacent, cosecant it hypotenuse over opposite

OpenStudy (anonymous):

hypotenuse is \[\sqrt{8^2+(3\sqrt{2})^2}\] \[=\sqrt{64+18}=\sqrt{82}\]